### v1

# Subtype-Aware Registration of Longitudinal Electronic Health Records

Abstract
Electronic Health Records (EHRs) contain extensive patient information that can inform downstream clinical decisions, such as mortality prediction, disease phenotyping, and disease onset prediction. A key challenge in EHR data analysis is the temporal gap between when a condition is first recorded and its actual onset time. Such timeline misalignment can lead to artificially distinct biomarker trends among patients with similar disease progression, undermining the reliability of downstream analysis and complicating tasks like disease subtyping. To address this challenge, we provide a subtype-aware timeline registration method that leverages data projection and discrete optimization to simultaneously correct timeline misalignment and improve disease subtyping. Through simulation and real-world data analyses, we demonstrate that the proposed method effectively aligns distorted observed records with the true disease progression patterns, enhancing subtyping clarity and improving performance in downstream clinical analyses.
footnotetext: 1Equal contribution. 2Corresponding author. E-mail: anru.zhang@duke.edu
1 Introduction
Electronic Health Records (EHR) are digitized collections of patient hospital information, encompassing data such as demographic, lab measurements, diagnoses, and medications. Data collected from intensive care unit (ICU) patients are of particular interest in clinical research due to their high frequency and timely collection. These data contain crucial information for executing timely disease interventions, managing patient care, and facilitating various downstream clinical decisions [1, 2, 3].
However, a significant challenge in utilizing EHR data lies in the temporal gap between when a condition is first recorded and its actual onset time. For example, a patient with pneumonia might be admitted to the hospital at the congestion stage (stage one) [4], while another patient might delay hospital visitation until a later stage due to factors like insurance coverage or personal health beliefs [5], resulting in unobserved data for the initial stages of disease progression. Conventional analyses on EHR data often overlook this misalignment of timelines, potentially leading to biased modeling results. Therefore, it is crucial to address this issue by develping methods that can effectively align observed EHR timelines with the true underlying disease progression.
1.1 Registration in the Literature
This need for timeline alignment is closely related to the problem of curve registration in functional data analysis, which seeks to align curves so that key features are properly synchronized across subjects [6, 7, 8]. In engineering, this problem is often referred to as time warping [9, 10]. Early methods aligned curves to a single baseline, using approaches like self-modeling nonlinear regression [11, 12] or template selection through functional principal component analysis (FPCA) [13], e.g., [14]. Subsequent work has explored a variety of alignment strategies. For instance, [15] proposed equating curve “moments” to capture local and global features, while [16] introduced a similarity index to iteratively register functional data. Other studies have applied Fisher-Rao metric-based approaches [17, 18, 19], computationally efficient techniques tailored for exponential family functional data [20, 21, 22], and pairwise warping strategies [23]. These approaches have been applied to a range of applications, including house price modeling [24], accelerometric activity data [25], and speech analysis [26]. However, since EHR data are often highly heterogeneous, sparse, noisy, and irregular, traditional curve registration methods are often not directly applicable.
Furthermore, some studies have explored combining curve registration with clustering to account for the latent subgroups among subjects. [27] used an EM algorithm with B-splines, while [28] employed a Bayesian framework that accommodates flexible warping functions. Yet, these methods often rely on specific data distributions that may not align with the complexities of EHR datasets.
To the best of our knowledge, [29] made a first attempt to address EHR-specific registration. They proposed aligning EHR trajectories by estimating time shifts to a common disease progression template. They assumed that all patients with a disease of interest share a common disease progression template and proposed a registration method using alternating optimization to align observed EHR data to this shared template. Despite its novelty, this approach has several limitations: (1) the reliance on linear interpolation and extrapolation for imputing missing data can lead to inaccurate estimations, particularly in sparsely observed EHR datasets; (2) the high computational burden of timeline registration for imputed curves becomes impractical for large patient cohorts; and (3) the assumption of a single progression template fails to account for the significant heterogeneity in diseases, where patients may follow distinct progression patterns.
In summary, while various registration and clustering methods exist, they remain constrained by assumptions unsuitable for EHR data’s sporadic and heterogeneous nature. These challenges necessitate the development of a new timeline registration method that effectively address the distinct complexities of EHR datasets, including data sparsity, diverse disease subtypes, and large patient populations.
1.2 Our Contributions
To address the limitations mentioned above, this paper introduces an efficient and effective subtype-aware registration method tailored for electronic health records data. Our approach recognizes that patients can be grouped into subtypes, each potentially following a different disease progression pattern. Specifically, we consider an EHR dataset comprising $N$ subjects, where each subject $i$ has longitudinal data $\{(t_{ik},X_{i}(t_{ik}))\mid k=1,2,\ldots,T_{i}\}$. Here, $t_{ik}$ denotes the observation times within the domain of interest $[T_{\min},T_{\max}]$, $X_{i}(t_{ik})$ represents the corresponding observations, and $T_{i}$ is the number of observations for subject $i$. The observation times and the number of observations vary significantly across subjects. See the left panel of Figure 1 for an illustrative example on trajectories. The subjects belong to multiple disjoint subtypes $\mathcal{P}_{1},\ldots,\mathcal{P}_{K}$, where each subtype $\mathcal{P}_{k}$ contains patients with similar disease progression patterns.
Our goal is to align the timelines of subjects within each subtype more closely to each other by estimating individual time shifts $\tau_{i}$ and clustering labels $\pi_{i}$ such that the shifted data $X_{i}(t-\tau_{i})$ align optimally for all $i$ in subtype $\mathcal{P}_{k}$. Recognizing that outliers commonly occur in the electronic health records [30], we also aim to ensure robustness against them in our procedure.
To achieve this, we transform the observed longitudinal data and all possible time shifts using cubic B-spline regression, reducing the data’s dimensionality and enabling accelerated computation. We then cluster the data in the transformed space and iteratively update the clusters by selecting the time shifts that minimize the distance to cluster centroids. Each cluster in the transformed space is considered a potential subtype of the disease of interest. By focusing on minimizing intra-cluster variance, our method effectively and robustly captures the heterogeneity of disease progression patterns while maintaining robustness to outliers. In summary, our proposed method overcomes the limitations of existing approaches by incorporating subtype awareness, improving computational efficiency, and enhancing robustness to data sparsity and outliers. We present a simulation example in Figure 1 to demonstrate the effectiveness of timeline registration. The left panel shows the observed data and the right panel shows the unobserved underlying patterns, which include two unknown subtypes of trajectories marked in blue and red. The recovered trajectories after applying the proposed timeline registration are displayed in the center panel of Figure 1. Further discussions on the simulation studies are deferred to Section 3.
Figure 1: Simulation 1 Trajectories: Left: Observed trajectories with unregistered timelines and unknown subtypes. Right: True trajectories with subtypes indicated by colors. Middle: Trajectories after applying the proposed timeline registration, with recovered subtypes marked by colors.
We evaluate our algorithm through extensive numerical studies. In six simulated datasets, our algorithm successfully realigned the observed data to the ground truth with less than one day of error in over 90% subjects. This realignment resulted in an enhanced separation between patient clusters, as measured by the Silhouette coefficient, from both data-driven and clinical perspectives. Furthermore, the realigned EHR data improved the predictive accuracy for severe AKI patients by 3–5%, demonstrating the effectiveness of the registered data in supporting accurate prediction modeling.
(a) Step 1: Time Series Transformation Using Cubic B-spline Regression
(b) Step 2: Iterations
(c) Step 3: Finalization
Figure 2: This figure illustrates subtype-aware registration using longitudinal EHR data. Black lines and dots represent the current curves, observations, and their low-dimensional embeddings. Red lines and dots indicate curves and observations that include potential time shifts under consideration. Blue dots represent potential outliers that were excluded from centroid evaluation for the time being.
2 Method for Subtype-Aware Registration of EHR
[…]
Next, we propose the EHR registration algorithm shown in Figure 2, which consists of three main steps, detailed in the following subsections. The pseudo-code for each step is provided in Algorithms 1, 2, and 3.
Step 1: Time Series Transformation Using Cubic B-spline Regression
In this step, we apply cubic B-spline regression [31, 32] to perform time series transformation. This technique models longitudinal data with flexible and smooth approximations, yielding a low-dimensional representation.
Specifically, we introduce $q$ internal knots, denoted as $\{\tau_{1},\ldots,\tau_{q}\}$, within the overall time domain $[T_{\min},T_{\max}]$. Based on these knots, we construct $p=q+3$ basis functions for cubic splines $B_{j,3}(t)$ recursively as follows.
First, the zeroth-degree basis functions $B_{j,0}(t)$ are defined as piecewise constants:
$B_{j,0}(t)=\left\{\begin{array}[]{ll}1&\text{if }t\in[\tau_{j},\tau_{j+1}),\\ 0&\text{otherwise}.\end{array}\right.$
Then, for $m=1,2,3$, the higher-degree basis functions are constructed recursively:
$B_{j,m}(t)=\frac{t-\tau_{j}}{\tau_{j+m}-\tau_{j}}B_{j,m-1}(t)+\frac{\tau_{j+m+1}-t}{\tau_{j+m+1}-\tau_{j+1}}B_{j+1,m-1}(t).$
For more details about B-splines, readers are referred to [33].
We perform B-spline regression using ridge regression (which provides more robustness) to fit both the original and time-shifted observations for each subject $i=1,\ldots,N$. Specifically, we pre-specify a set of potential shift positions $\Lambda=\{l_{0},l_{1},\ldots,l_{m}\}$, where $l_{0}=0$ means no shift. For each $i=1,\ldots,N$ and $l\in\Lambda$, we perform ridge regression $X_{i}(t-l)\sim B_{j,3}(t)$ by evaluating the following estimator:
$\widehat{\boldsymbol{\beta}}_{i,l}=\mathop{\rm arg\min}_{\boldsymbol{\beta}_{i,l}}\left\{\sum_{\begin{subarray}{c}k=1\\ T_{\min}\leq t_{ik}-l\leq T_{\max}\end{subarray}}^{T_{i}}\left(X_{i}(t_{ik}-l)-\sum_{j=1}^{p}\beta_{ij}B_{j,3}(t_{ik})\right)^{2}+\lambda\sum_{j=1}^{p}\beta_{ij}^{2}\right\},$
where $\lambda$ is the tuning parameter that controls the penalization level, and $\widehat{\boldsymbol{\beta}}_{i,l}=(\widehat{\beta}_{i1},\widehat{\beta}_{i2},\ldots,\widehat{\beta}_{ip})$ is the vector of estimated coefficients. In this way, the underlying structure and dynamics of the longitudinal data are reduced to $p$ coefficients. This cubic B-spline regression yields the following low-dimensional embedding:
$\{(t_{ik},X_{i}(t_{ik}))\mid k=1,2,\ldots,T_{i}\}\quad\longrightarrow\quad\widehat{\boldsymbol{\beta}}_{i,l}.$
We compute and store these low-dimensional representations for all subjects and potential shifts into a tensor $\mathbf{\Omega}\in\mathbb{R}^{N\times|\Lambda|\times p}$:
$\mathbf{\Omega}_{[i,l,:]}=\widehat{\boldsymbol{\beta}}_{i,l},\quad\text{for }i=1,\ldots,N;\quad l\in\Lambda.$
Step 2: Iterations
We set the initial time shifts as $\delta_{i}^{(0)}=0$ for all $i=1,\ldots,N$. Next, we run the following steps for $t=1,2,\ldots$:
  1. 1.
    Clustering with Optimal Number of Clusters. For each candidate number of clusters $k\in\{2,\ldots,M\}$, we conduct clustering (e.g., $k$-means, $k$-medoids) with $k$ clusters on the low-dimensional representations with time shifts $\delta_{i}^{(t-1)}$: $\{\mathbf{\Omega}_{i,\delta_{i}^{(t-1)}}\}_{i=1}^{N}.$ When $t=1$, the data is the original longitudinal dataset without time shifts.
    We determine the number of clusters $K^{(t)}$ by selecting the $k$ that maximizes the average Silhouette coefficient [34]:
    $K^{(t)}=\mathop{\rm arg\max}_{k\in\{2,\ldots,M\}}s(k),\quad s(k)=\frac{1}{N}\sum_{i=1}^{N}s(i,k),$ (1)
    where $s(i,k)$ is the Silhouette coefficient for subject $i$ when the number of clusters is $k$.
    This results in $K^{(t)}$ clusters of the subjects: $\{\mathcal{C}_{1}^{(t)},\ldots,\mathcal{C}_{K^{(t)}}^{(t)}\}\subseteq[N]$, where $\mathcal{C}_{i}^{(t)}\cap\mathcal{C}_{j}^{(t)}=\emptyset$ for all $i\neq j$.
  2. 2.
    Iterative Updates. For each cluster $k=1,\ldots,K^{(t)}$, we perform:
    • Compute the centroid of cluster $\mathcal{C}_{k}^{(t)}$:
      $\bar{\boldsymbol{\beta}}_{k}^{(t)}=\frac{1}{|\mathcal{C}_{k}^{(t)}|}\sum_{i\in\mathcal{C}_{k}^{(t)}}\mathbf{\Omega}_{i,\delta_{i}^{(t-1)}}.$
    • Calculate the Euclidean distance between each point in the cluster and the centroid:
      $d_{i}=\left\|\mathbf{\Omega}_{i,\delta_{i}^{(t-1)}}-\bar{\boldsymbol{\beta}}_{k}^{(t)}\right\|,\quad\forall i\in\mathcal{C}_{k}^{(t)}.$
    • Select the subset $\mathcal{S}_{k}^{(t)}$ of points corresponding to the smallest $\left\lceil\alpha|\mathcal{C}_{k}^{(t)}|\right\rceil$ distances:
      $\mathcal{S}_{k}^{(t)}=\left\{i\in\mathcal{C}_{k}^{(t)}\mid d_{i}\text{ is among the smallest }\left\lceil\alpha|\mathcal{C}_{k}^{(t)}|\right\rceil\text{values}\right\},$
      where $0<\alpha\leq 1$ is a predefined proportion.
    • Recompute the centroid using only the selected points:
      $\bar{\boldsymbol{\gamma}}_{k}^{(t)}=\frac{1}{|\mathcal{S}_{k}^{(t)}|}\sum_{i\in\mathcal{S}_{k}^{(t)}}\mathbf{\Omega}_{i,\delta_{i}^{(t-1)}}.$
      This step removes potential outliers ($\mathcal{C}_{k}^{(t)}\setminus\mathcal{S}_{k}^{(t)}$) from the centroid computation, enhancing robustness.
    • Update the time shifts for each subject $i\in\mathcal{C}_{k}^{(t)}$ by finding the shift $\delta_{i}^{(t)}\in\Lambda$ that minimizes the distance to the updated centroid:
      $\delta_{i}^{(t)}=\mathop{\rm arg\min}_{\delta\in\Lambda}\left\|\mathbf{\Omega}_{i,\delta}-\bar{\boldsymbol{\gamma}}_{k}^{(t)}\right\|^{2}.$
  3. 3.
    Stopping Criterion. We employ stopping criteria to ensure efficient convergence and prevent excessive computation. Based on the initial clustering $\{\mathcal{C}_{1}^{(1)},\ldots,\mathcal{C}_{K^{(1)}}^{(1)}\}$, we evaluate the average Silhouette coefficient $s(K^{(1)})$ using Equation (1). The maximum number of iterations $I$ is set as:
    $I=\left\{\begin{array}[]{ll}I_{\text{small}}&\text{if }s(K^{(1)})>\tau,\\ I_{\text{large}}&\text{if }s(K^{(1)})\leq\tau,\end{array}\right.$
    where $\tau$ is a predefined threshold.
    We terminate the iterations when:
    $\left\{K^{(t)}=K^{(t-1)}\text{ and }t>I\right\}\quad\text{or}\quad\left\{t>2I\right\}.$
    The condition $t>2I$ provides a safeguard to ensure termination even if convergence is not achieved within the expected number of iterations.
2.1 Step 3: Finalization
Upon termination at iteration $t=t_{\text{final}}$, we address the data points not selected during the last iteration. For each cluster $k$, consider the set of unselected points: $\mathcal{C}_{k}^{(t_{\text{final}})}\setminus\mathcal{S}_{k}^{(t_{\text{final}})}$, where $\mathcal{S}_{k}^{(t_{\text{final}})}$ denotes the subset of selected points within cluster $k$ after the final iteration.
For each unselected point $j\in\mathcal{C}_{k}^{(t_{\text{final}})}\setminus\mathcal{S}_{k}^{(t_{\text{final}})}$, find the optimal shift $\delta_{j}^{*}$ and assign it to the most appropriate cluster center by solving:
$\delta_{j}^{*},\ k^{\prime*}=\mathop{\rm arg\min}_{\delta_{j}\in\Lambda,\ k^{\prime}=1,\ldots,K^{(t_{\text{final}})}}\left\|\mathbf{\Omega}_{j,\delta_{j}}-\bar{\boldsymbol{\gamma}}_{k^{\prime}}^{(t_{\text{final}})}\right\|.$
This assigns $j$ to $\mathcal{S}_{k}^{(t_{\text{f}inal})}$ and shift state that best represents it.
The final output of the algorithm is the number of clusters $\widehat{K}^{t_{\rm final}}$, the set of shift states, and clusters for all subjects:
$\{\delta_{i}^{*}:i=1,2,\ldots,N\}\quad\text{and}\quad\{\mathcal{S}_{k}^{(t_{\text{final}})}\}_{k\in\widehat{K}^{(t_{\text{final}})}}.$
The overall pseudo code is summarized in Algorithms 1-3.
Algorithm 1 Time Series Clustering with Dynamic Time Shifts: Step 1: Transformation
1:Longitudinal data $\{(t_{ik},X_{i}(t_{ik}))\mid i=1,\ldots,N;\ k=1,\ldots,T_{i}\}$, Potential shift positions $\Lambda=\{l_{0},l_{1},\ldots,l_{m}\}$, Maximum number of clusters $M$, predefined threshold $\tau$, Parameters $\alpha$, $I_{\text{small}}$, $I_{\text{large}}$, penalization parameter $\lambda$.
2:Optimal shift states $\{\delta_{i}^{*}\mid i=1,2,\ldots,N\}$ and clusters $\{\mathcal{C}_{k}^{(t_{\text{final}})}\}$
3:Introduce internal knots $\{\tau_{1},\ldots,\tau_{q}\}$ in $[T_{\min},T_{\max}]$
4:Construct cubic B-spline basis functions $B_{j,3}(t)$
5:for each subject $i=1$ to $N$ do
6:     for each shift $l\in\Lambda$ do
7:         Perform ridge regression to obtain coefficients:
$\widehat{\boldsymbol{\beta}}_{i,l}=\mathop{\rm arg\min}_{\boldsymbol{\beta}_{i,l}}\left\{\sum_{\begin{subarray}{c}k=1\\ T_{\min}\leq t_{ik}-l\leq T_{\max}\end{subarray}}^{T_{i}}\left(X_{i}(t_{ik}-l)-\sum_{j=1}^{p}\beta_{ij}B_{j,3}(t_{ik})\right)^{2}+\lambda\sum_{j=1}^{p}\beta_{ij}^{2}\right\}$
8:         Store coefficients in tensor $\mathbf{\Omega}_{[i,l,:]}=\widehat{\boldsymbol{\beta}}_{i,l}$
9:     end for
10:end for
Algorithm 2 Step 2: Iterative Clustering and Time Shift Adjustment
1:Initialize time shifts $\delta_{i}^{(0)}=0$ for all $i$. Set iteration counter $t\leftarrow 1$
2:repeat
3:     for each $k$ in $\{2,\ldots,M\}$ do
4:         Perform clustering with $k$ clusters on $\{\mathbf{\Omega}_{i,\delta_{i}^{(t-1)}}\}_{i=1}^{N}$
5:         Compute average Silhouette coefficient $s(k)$ using (1)
6:     end for
7:     Select optimal number of clusters: $K^{(t)}=\mathop{\rm arg\max}_{k}s(k)$; obtain clusters $\{\mathcal{C}_{1}^{(t)},\ldots,\mathcal{C}_{K^{(t)}}^{(t)}\}$
8:     for each cluster $k=1$ to $K^{(t)}$ do
9:         Compute centroid: $\bar{\boldsymbol{\beta}}_{k}^{(t)}=\sum_{i\in\mathcal{C}_{k}^{(t)}}\mathbf{\Omega}_{i,\delta_{i}^{(t-1)}}/|\mathcal{C}_{k}^{(t)}|$
10:         Calculate distances $d_{i}=\left\|\mathbf{\Omega}_{i,\delta_{i}^{(t-1)}}-\bar{\boldsymbol{\beta}}_{k}^{(t)}\right\|$ for $i\in\mathcal{C}_{k}^{(t)}$
11:         Select subset $\mathcal{S}_{k}^{(t)}$ of points with smallest $\left\lceil\alpha|\mathcal{C}_{k}^{(t)}|\right\rceil$ distances
12:         Recompute centroid: $\bar{\boldsymbol{\gamma}}_{k}^{(t)}=\sum_{i\in\mathcal{S}_{k}^{(t)}}\mathbf{\Omega}_{i,\delta_{i}^{(t-1)}}/\frac{1}{|\mathcal{S}_{k}^{(t)}|}$
13:     end for
14:     for each subject $i=1$ to $N$ do
15:         Update time shift: $\delta_{i}^{(t)}=\mathop{\rm arg\min}_{\delta\in\Lambda}\left\|\mathbf{\Omega}_{i,\delta}-\bar{\boldsymbol{\gamma}}_{k}^{(t)}\right\|^{2}$, where $k$ is the cluster containing subject $i$
16:     end for
17:     Update $t\leftarrow t+1$
18:     Check Stopping Criteria
19:     if $K^{(t)}=K^{(t-1)}$ and $t>I$ then
20:         Terminate Iteration
21:     else if $t>2I$ then
22:         Terminate Iteration (Hard Stop)
23:     end if
24:until Termination condition is met
Algorithm 3 Step 3: Iterative Clustering and Time Shift Adjustment
1:for each cluster $k$ do
2:     for each unselected point $j\in\mathcal{C}_{k}^{(t_{\text{final}})}\setminus\mathcal{S}_{k}^{(t_{\text{final}})}$ do
3:         Assign optimal shift and cluster by evaluating
$\delta_{j}^{*},\ k^{\prime*}=\mathop{\rm arg\min}_{\delta_{j}\in\Lambda,\ k^{\prime}=1,\ldots,K^{(t_{\text{final}})}}\left\|\mathbf{\Omega}_{j,\delta_{j}}-\bar{\boldsymbol{\gamma}}_{k^{\prime}}^{(t_{\text{final}})}\right\|.$
Then reassign $j$ to cluster $\mathcal{S}_{k^{\prime}*}^{(t_{\text{final}})}$.
4:     end for
5:end for
6:Output final shift amounts $\{\delta_{i}^{*}\mid i=1,2,\ldots,N\}$ and clusters $\{\mathcal{S}_{k}^{(t_{\text{final}})}\}$
3 Simulations
In this section, we conduct simulation studies across six scenarios to evaluate the effectiveness of our proposed algorithm. These scenarios covered a variety of different longitudinal data structures, including different shapes, different means, and their combinations. The objective is to assess the proposed algorithm’s ability to correctly recover the true temporal positions of these longitudinal data under various conditions. Next, we will describe the common setup, then discuss specific designs, and finally discuss the simulation results.
Common Simulation Setup.
Data Generation: For each subject $i$ (where $i=1,2,\ldots,1000$), the observation time points $t_{ij}$ (where $j=1,2,\ldots,28$) were randomly drawn from a continuous uniform distribution over the interval $[1,21]$ ($t_{ij}\sim\text{Uniform}(1,21)$). These time points were then sorted in ascending order for each subject. In all scenarios, the noise terms $\epsilon_{ij}$ were randomly drawn from a normal distribution $\epsilon_{ij}\sim N(0,0.8^{2})$.
Temporal Misalignment Simulation: In all scenarios, each subject was assigned a time shift to simulate the delay between the onset of illness and hospital admission. Specifically, in each scenario, 60% of the subjects in a group had a time shift of 0, representing patients who were admitted to the hospital immediately upon illness onset or were promptly transferred from general wards to the ICU after the onset of severe symptoms. The remaining 40% of subjects were assigned a random time shift ranging from 1 to 4 days (integers), simulating a delay of 1 to 4 days between illness onset and hospital admission. After applying the time shifts, any observations with time points less than 1 were removed to simulate the missing data during the “onset-to-admission” period when patients experience a delay in hospital admission. Additionally, observations with time points greater than 17 were removed to ensure a consistent study time window across all individuals.
Algorithm Parameter Settings: We fix 2 internal knots at 8 and 13. The potential shift positions $\Lambda$ were set to $\{0,1,2,3,4\}$. The range of cluster numbers was set to $\{2,3,\ldots,8\}$, i.e., $M=8$. The selection proportion $\alpha$ was set to 0.95, the threshold $\tau$ was set to 0.45, and the maximum iteration numbers were set to $I_{\text{small}}=1$ and $I_{\text{large}}=10$. We employ $k$-medoids as the clustering method employed given its robustness to outliers [35].
Scenario 1:
This scenario involves two types of trajectories that are similar in shape but have significant differences in mean, defined as follows:
$\begin{split}\text{Group 1 (500 subjects): }\quad&Y_{i}^{(1)}(t_{ij})=20+3\sin(0.6(t_{ij}+4))+\epsilon_{ij};\\ \text{Group 2 (500 subjects): }\quad&Y_{i}^{(2)}(t_{ij})=17+3\sin(0.6(t_{ij}+4))+\epsilon_{ij}.\end{split}$
In this setting, the trajectories from Group 1 had a higher mean level compared to Group 2, while the overall shape of their trajectories, dictated by the sinusoidal base sequence, remained similar. This setup allows us to evaluate the algorithm’s ability to distinguish between groups with different mean levels but similar temporal patterns.
Scenario 2:
This scenario involves two types of trajectories with distinct shape differences but similar overall means.
$\begin{split}\text{Group 1 (500 subjects): }\quad&Y_{i}^{(1)}(t_{ij})=17+2\sin(0.6(t_{ij}+4))+\epsilon_{ij};\\ \text{Group 2 (500 subjects): }\quad&Y_{i}^{(2)}(t_{ij})=2t_{ij}+3\sin(0.6(t_{ij}+4))+\epsilon_{ij}.\end{split}$
While both groups have similar overall means, Group 1 follows a sinusoidal pattern, whereas Group 2 exhibits a combination of a sinusoidal pattern and a linear trend.
Scenario 3:
This scenario involves three types of trajectories defined as follows:
$\begin{split}\text{Group 1 (300 subjects): }\quad&Y_{i}^{(1)}(t_{ij})=20+3\sin(0.6(t_{ij}+4))+\epsilon_{ij};\\ \text{Group 2 (400 subjects): }\quad&Y_{i}^{(2)}(t_{ij})=17+3\sin(0.6(t_{ij}+4))+\epsilon_{ij};\\ \text{Group 3 (300 subjects): }\quad&Y_{i}^{(3)}(t_{ij})=16+0.5t_{ij}+3\sin(0.9(t_{ij}+4))+\epsilon_{ij}.\end{split}$
In summary, Groups 1 and 2 have similar shapes but different means, while Group 3 has a different shape that overlaps with the first two groups. This scenario helps evaluate the algorithm’s ability to distinguish between groups with both shape and mean differences.
Scenario 4:
This scenario involves four types of trajectories with shape differences and overlaps:
$\begin{split}\text{Group 1 (250 subjects):}\quad&Y_{i}^{(1)}(t_{ij})=24+3\sin(0.6(t_{ij}+4))+\epsilon_{ij};\\ \text{Group 2 (250 subjects):}\quad&Y_{i}^{(2)}(t_{ij})=17+3\sin(0.6(t_{ij}+4))+\epsilon_{ij};\\ \text{Group 3 (250 subjects):}\quad&Y_{i}^{(3)}(t_{ij})=38-0.5t_{ij}+4\sin(0.9(t_{ij}+3))+\epsilon_{ij};\\ \text{Group 4 (250 subjects):}\quad&Y_{i}^{(4)}(t_{ij})=21+0.9t_{ij}+3\sin(0.6(t_{ij}+4))+\epsilon_{ij}.\end{split}$
In this scenario, Groups 1 and 2 have similar shapes but different means, while Groups 3 and 4 exhibit different shapes that overlap with each other and the first two groups.
Scenario 5:
This scenario includes 4 groups of trajectories and each group includes 250 subjects. The trajectories of each group is generated as follows:
$\begin{split}\text{Group 1 (250 subjects):}\quad Y_{i}^{(1)}(t_{ij})=&24+3\sin(0.6(t_{ij}+4))+\epsilon_{ij}^{(1)};\\ \text{Group 2 (250 subjects):}\quad Y_{i}^{(2)}(t_{ij})=&17+3\sin(0.6(t_{ij}+4))+\epsilon_{ij}^{(2)};\\ \text{Group 3 (250 subjects):}\quad Y_{i}^{(3)}(t_{ij})=&28-0.5t_{ij}+4\sin(0.9(t+3))+\epsilon_{ij}^{(3)};\\ \text{Group 4 (250 subjects):}\quad Y_{i}^{(4)}(t_{ij})=&21-0.5t_{ij}+4\sin(0.9(t+3))+\epsilon_{ij}^{(4)}.\end{split}$
In this scenario, the trajectories in Groups 1 and 2 share one shape with different means, while Groups 3 and 4 share another shape with different means. This scenario helps evaluate the algorithm’s ability to distinguish between overlapping groups with similar shapes but different mean levels.
Scenario 6:
This scenario includes 4 groups of trajectories and each group includes 250 subjects. The trajectories of each group is generated as follows:
$\begin{split}\text{Group 1 (250 subjects):}\quad&Y_{i}^{(1)}(t_{ij})=23+3\sin(0.6(t_{ij}+4))+\epsilon_{ij};\\ \text{Group 1 (250 subjects):}\quad&Y_{i}^{(2)}(t_{ij})=17+3\sin(0.6(t_{ij}+4))+\epsilon_{ij};\\ \text{Group 1 (250 subjects):}\quad&Y_{i}^{(3)}(t_{ij})=20+3\sin(0.6(t_{ij}+4))+\epsilon_{ij};\\ \text{Group 1 (250 subjects):}\quad&Y_{i}^{(4)}(t_{ij})=25-0.5t_{ij}+4\sin(0.9(t_{ij}+3))+\epsilon_{ij}.\end{split}$
In summary, Groups 1, 2, and 3 share a similar shape with different means, while Group 4 has a distinctly different shape but overlaps with other types. This scenario helps evaluate the algorithm’s ability to distinguish between groups with both shape and mean differences, including overlapping patterns.
Scenario Rate of Exact Recovery Rate of Recovery Error $\leq$ 1 MAE (days) Runtime (min)
1 (2 clusters) 69.0% [62.0%, 83.9%] 92.0% [84.9%, 99.5%] 0.40 [0.17, 0.55] 3.13 [1.22, 7.42]
2 (2 clusters) 75.8% [73.1%, 80.0%] 94.5% [90.6%, 98.0%] 0.35 [0.24, 0.49] 1.24 [0.86, 1.93]
3 (3 clusters) 74.0% [71.1%, 78.1%] 91.5% [90.5%, 93.1%] 0.40 [0.34, 0.43] 1.25 [0.87, 1.93]
4 (4 clusters) 69.2% [61.3%, 79.6%] 97.1% [95.7%, 98.5%] 0.37 [0.28, 0.47] 2.63 [1.86, 5.46]
5 (4 clusters) 65.2% [50.7%, 75.9%] 90.6% [87.9%, 93.2%] 0.60 [0.45, 0.77] 4.56 [1.29, 7.57]
6 (4 clusters) 76.0% [69.8%, 80.7%] 93.9% [89.8%, 96.4%] 0.41 [0.30, 0.58] 1.69 [1.14, 2.07]
Table 1: Simulation Results: We present the average value and 95% confidence interval for the rate of exact recovery, rate of recovery error, mean absolute value, and runtime
We use four metrics to evaluate the performance of the algorithm: Accuracy Rate refers to the proportion of subjects for whom the algorithm successfully recovered the correct temporal positions of their longitudinal data; Probability of Error $\leq 1$ Day refers to the proportion of subjects with a result error of less than or equal to 1 day; Mean Absolute Error (MAE) represents the average time difference between the recovered temporal positions and the correct temporal positions for all subjects, and the Runtime indicates the time taken for the algorithm to run once on the simulated dataset. The simulation results, which present the average values and 95% confidence intervals for all four metrics over 500 runs, are shown in Table 1. The visualizations of the simulation results can be found in Figures 1, 3, 4, 5, 6, and 7 in the Appendix.
In summary, our algorithm effectively recovered the temporal positions from the simulated data across all six scenarios. The evaluation metrics show similar values regardless of the complexity of the scenario, indicating the robustness of our algorithm. The total runtime per execution of the algorithm on the simulated dataset does not exceed 2 minutes, demonstrating its low computational complexity and potential for practical application on large-scale EHR datasets.
4 Comprehensive Analyses of Longitudinal EHR Data in Acute Kidney Injury Patients
Acute Kidney Injury (AKI) is a common clinical condition characterized by an abrupt decline in kidney function, resulting from various causes such as decreased renal perfusion, direct nephron injury, or urinary tract obstruction [36]. AKI can develop rapidly within hours or days and is associated with significant morbidity and high mortality rates [37]. This syndrome is highly heterogeneous, and the complexity of its pathophysiology poses significant challenges in developing effective patient treatment plans [37].
In this study, we utilize longitudinal EHR data from AKI patients to demonstrate that our proposed method effectively aligns observed EHR records among different subjects who belong to the same cluster. Specifically, we employ unsupervised learning techniques, complemented by clinical evaluations and supervised learning, to validate the effectiveness of our subtype-aware EHR registration method.
4.1 Experimental Setup
Dataset and Cohort Selection
We conduct our real data analysis using the Medical Information Mart for Intensive Care (MIMIC)-IV database [38, 39]. The MIMIC-IV dataset contains comprehensive information on over 40,000 patients admitted to intensive care units (ICUs) at the Beth Israel Deaconess Medical Center (BIDMC) between 2008 and 2019. This dataset includes a wide range of clinical data such as demographics, vital signs, laboratory results, procedures, diagnoses, and medications.
To select our cohort of patients with Acute Kidney Injury (AKI), we identified ICU stays where patients had ICD (International Classification of Diseases) codes linked to acute kidney failure, ensuring that the accompanying text explanations contained the keyword “acute kidney injury.” We excluded patients who, during their ICU stay, were labeled with burns, had end-stage renal disease, or renal dialysis. Additionally, patients with an estimated glomerular filtration rate (eGFR) of less than 15 mL/min/1.73 m2 were excluded to avoid including those with end-stage renal disease [40].
For data quality assurance, we removed ICU stays with fewer than 15 recorded measurements within the first 21 days after ICU admission. Furthermore, we excluded ICU stays that had two or fewer recorded measurements in each of the following intervals since ICU admission: 0–7 days, 7–14 days, and 14–21 days. The key feature used in our analysis was serum creatinine levels, a laboratory measurement that is the primary determinant for diagnosing and monitoring AKI [41].
Data Preprocessing and Model Selection
We perform unsupervised learning on the data. Specifically, we transform the original longitudinal data as detailed in Section 2, and used the estimated coefficients of the spline basis functions to represent each ICU stay. We applied $k$-means clustering [42] and $k$-medoids clustering [43] on these representations. The quality of the resulting clusters was evaluated by computing the average Silhouette coefficient [34].
For clinical evaluation, we labeled each ICU stay based on the AKI severity stages defined below [41, 44]:
  1. 1.
    Stage 1: Serum creatinine (SCr) level rises to 1.5×1.5$\times$1.5 × to 1.9×1.9$\times$1.9 × the baseline value within seven days since time 0.
  2. 2.
    Stage 2: SCr level rises to 2.0×2.0$\times$2.0 × to 2.9×2.9$\times$2.9 × the baseline value within seven days since time 0.
  3. 3.
[…]
    1. (a)
[…]
    2. (b)
      Subtype 2: SCr level rises to 3×3$\times$3 × or greater than the baseline value within seven days since time 0 and does not meet the criteria for Subtype 1.
The baseline SCr value was computed using the Modification of Diet in Renal Disease (MDRD) equation [45].
[…]
4.2 Unsupervised Learning – Phenotyping
We first evaluate the impact of timeline registration on clustering AKI ICU stays using $k$-medoids and $k$-means. The average Silhouette coefficients for different number of clusters $K\in[2,8]$ are presented in Table 2. Notably, when $K=2$ and $K=3$, the clustering results yield significantly higher average Silhouette coefficients, indicating more distinct clustering results. Additionally, we observe that both $k$-medoids and $k$-means show substantial increases in the average Silhouette coefficients after applying the proposed timeline registration method. Given the heterogeneity of AKI, we expect the ICU stays to cluster according to different subtypes of the condition. The numerical results demonstrate the effectiveness of our timeline registration method in aligning the timelines of patients within the same subtype.
$k$-medoids $k$-means
Observed Data Registered Data Registered Data Registered Data
K=2 0.431 0.483 0.584 0.628
K=3 0.226 0.260 0.523 0.545
K=4 0.249 0.269 0.325 0.332
K=5 0.208 0.280 0.350 0.380
K=6 0.208 0.271 0.352 0.395
K=7 0.111 0.112 0.346 0.402
K=8 0.108 0.159 0.284 0.220
Table 2: Comparison of average Silhouette coefficients for original and adjusted data after applying the proposed timeline registration method, using $k$-medoids and $k$-means clustering.
In computational phenotyping research, the number of subtypes $K$ is often unknown and must be determined. Therefore, we consider a scenario in which the number of clusters falls within a pre-specified range, and we apply the proposed timeline registration method to determine this number in a data-driven manner. In Table 3, we present results for two scenarios: one requiring a smaller number of clusters and the other requiring a larger number, as potentially suggested by auxiliary information. In both cases, we observe that data after timeline registration show increased average Silhouette coefficients, indicating that the timeline registration method provides consistent improvement in downstream clustering performance.
$k$-medoids $k$-means
Observed Data Registered Data Observed Data Registered Data
$K\in[2,4]$ 0.431 0.483 0.584 0.628
$K\in[5,8]$ 0.208 0.283 0.352 0.431
Table 3: Comparison of average Silhouette coefficients for original and adjusted data after applying the proposed timeline registration method, using $k$-medoids and $k$-means clustering. Number of clusters $K$ is either in the range of $[2,4]$ or $[5,8]$.
4.3 Clinical Evaluation
Next, we incorporate clinical information, specifically the AKI severity stages, into the evaluation of timeline registration performance. Recall that the definition of AKI severity stages, as outlined in Section 4.1, relies on the starting time point. This implies that the AKI severity stage of an individual subject may change after applying timeline registration. Therefore, for both the observed data (without timeline registration) and the data after applying the proposed timeline registration procedure, we label each subject as Stage 1 (S1), Stage 2 (S2), Stage 3 Subtype 1 (S3-1), or Stage 3 Subtype 2 (S3-2). We then group both the observed and registered datasets into five different configurations, as described in the first column of Table 4, and evaluate the average Silhouette coefficient for each configuration. The results for these five different AKI sub-category divisions are presented in Table 4. Consistently, the data processed through timeline registration demonstrate an increased average Silhouette coefficient, indicating improved data aggregation within each AKI sub-category.
Observed Data Registered Data
4-Groups (S1, S2, S3-1, S3-2) -0.049 -0.041
3-Groups (S1, S2, (S3-1, S3-2)) -0.057 -0.047
3-Groups ((S1, S2), S3-1, S3-2) 0.286 0.327
2-Groups ((S1, S2), (S3-1, S3-2)) 0.321 0.341
2-Groups ((S1, S2, S3-1), S3-2 ) 0.527 0.627
Table 4: Comparison of observed and recovered data after registration in the clinical setting using new criteria for defining creatinine baseline level.
4.4 Supervised Learning – Severe AKI Prediction
To validate the effectiveness of timeline registration in enhancing downstream clinical analysis, we perform severe AKI [^1] See definition of severe AKI in Data Preprocessing and Model Section in Section 4.1. prediction and compare the performance using observed data and the recovered data after registration. The experiment was conducted 500 times and we report the mean value with its 95% confidence interval, presented in Table 5 and Table 6.
Observed Data Registered Data ($k$-medoids) Registered Data ($k$-means)
AUROC 0.8050 [0.7836, 0.8272] 0.8504 [0.8316, 0.8688] 0.8456 [0.8242, 0.8658]
Accuracy 0.8164 [0.8015, 0.8309] 0.8343 [0.8207, 0.8487] 0.8278 [0.8138, 0.8419]
AUPRC 0.8023 [0.7812, 0.8237] 0.8379 [0.8199, 0.8572] 0.8289 [0.8081, 0.8499]
Table 5: Comparisons of severe AKI prediction using 3-day observed data and recovered data after registration. Metrics are presented using mean values with 95% confidence intervals.
From both tables, we observe that using data after registration results in significant increases in the AUROC (area under the receiver operating characteristic curve) and AUPRC (area under the precision-recall curve) values. Accuracy values also slightly increase. This comparison indicates that timeline registration is influential in enhancing downstream clinical analysis and can be considered an essential preprocessing step.
Observed Data Registered Data ($k$-medoids) Registered Data ($k$-means)
AUROC 0.8276 [0.8062, 0.8489] 0.8787 [0.8621, 0.8957] 0.8776 [0.8587, 0.8972]
Precision 0.8138 [0.7894, 0.8374] 0.8318 [0.8116, 0.8548] 0.8251 [0.8035, 0.8449]
Recall 0.7326 [0.7104, 0.7543] 0.7684 [0.7451, 0.7903] 0.7581 [0.7372, 0.7807]
Accuracy 0.8273 [0.8125, 0.8412] 0.8463 [0.8323, 0.8604] 0.8403 [0.8261, 0.8546]
AUPRC 0.8256 [0.8046, 0.8460] 0.8680 [0.8509, 0.8861] 0.8621 [0.8433, 0.8806]
Table 6: Comparisons of severe AKI prediction using 4-day observed data and recovered data after registration. Metrics are presented using mean values with 95% confidence intervals.
5 Discussions and Conclusions
In this paper, we demonstrate the necessity of performing timeline registration for longitudinal EHR data, aligning patient records to reflect the intrinsic disease progression trends. To evaluate the feasibility of the proposed registration algorithm for recovering ground truth longitudinal data, we performed simulations with six different scenarios. The results indicate that the proposed method can effectively recover observed data to ground truth with low time complexity. To demonstrate the impact of registration on downstream clinical data analysis, we perform clustering with clinical evaluation and severe AKI prediction and compare their results using data with or without timeline registration. The results suggest that registration can effectively improve the performance of downstream clinical data analysis. We also conclude that the registration algorithm is of high stability, reflected by the minimal 95% confidence intervals given 500 times repetition of experiments. Additionally, our proposed registration algorithm has great generalizability in that its clustering step can be substituted by any desired clustering method. Registration on continuous space and evaluation of registration on other critical medical events will be investigated in our future work. EHR timeline registration is an emerging field and our work provides insight for improving clinical analysis using EHR data.